OpenStudy (anonymous):

a 50 kg carte is lifted to a height of 10 meters. how much potential energy does the crate have after being lifted?

OpenStudy (earthcitizen):

\[P=mgh\]

OpenStudy (earthcitizen):

P.E= (50)*(9.81)*(10)=4905J

OpenStudy (anonymous):

if it ook 5 seconds to lift the carte using a machine what was the power rating of the machine

OpenStudy (earthcitizen):

\[Power= delta W/delta t\]

OpenStudy (earthcitizen):

\[Power= \delta W/\delta t\]

OpenStudy (anonymous):

s tha answer should be

OpenStudy (earthcitizen):

plug in the work done and divide by the time taken

OpenStudy (earthcitizen):

Power = 4905/5=981J/s